# The Hidden Role of Anisotropies in Shaping Structure Formation in Cosmological N-Body Simulations

### v1

@@ III.1 Correlation function and power spectrum
We will determine the $n$-dimensional correlation function $\xi_{n\text{d}}(r)$ and power spectrum $P_{n\text{d}}(k)$, where $n=2,3$, for the distribution of particles in a simulation box of side length $L$, assuming periodic boundary conditions. For the three-dimensional case ($n=3$), we consider the full volume of the box. We also analyze quasi-two-dimensional slices, i.e., slabs with vertical thickness $\Delta z$ much smaller than the box side length, extending across both the $x$ and $y$ directions. Within each slice, we compute the two-dimensional correlation function $\xi_{2\text{d}}(r)$ and power spectrum $P_{2\text{d}}(k)$. Anisotropies introduced by the lattice used in the pre-initial conditions are still present even in these two-dimensional slices. Since the numerical analysis in two dimensions is simpler and faster than in three dimensions, we will primarily present results in 2D. We note that, for both simulations, the results presented correspond to a single two-dimensional slice. However, we have repeated the analysis across multiple independent slices, finding similar results in all cases. This confirms that the observed features are robust and not specific to a particular projection. As the statistical properties averaged over spherical (or circular) regions have been extensively studied in the literature, we do not repeat such an analysis in this work.
@@ III.2 Angular Distribution of Pairwise Distances
Estimations of the two-point correlation function or the power spectrum are typically based on the assumption that the underlying distribution is isotropic. However, it is precisely this assumption that we aim to test. To determine whether a point distribution is isotropic or exhibits preferred orientations, one can analyze spatial anisotropies using a range of statistical and geometric methods. We have focused on one particular statistical measure after testing several possibilities. This choice does not imply that it is necessarily the most suitable or optimal diagnostic; rather, it demonstrates that this statistic is effective in identifying anisotropic features in the point distribution.
@@
As mentioned above, given the geometry of the system, we focus on quasi-two-dimensional slices. Although structure formation is inherently a fully three-dimensional dynamical process, it is important to consider that particles typically move only a few Mpc over the entire time evolution. Consequently, from a statistical perspective, it is sufficient to isolate a two-dimensional slice of thickness greater than $\sim 5$ Mpc in order to track approximately the same subsample of particles throughout the evolution.
Moreover, the symmetry of the lattice used as a pre-initial condition justifies reducing the analysis of spatial anisotropies to a two-dimensional slice. This approach is not only numerically more efficient but also retains the essential features of directional correlations that are relevant to the problem.
@@
The generalization of the ADPD to quantify angular anisotropies in a three-dimensional (3D) point distribution, is straightforward. The direction of a separation vector is specified by two angles in spherical coordinates: the azimuthal angle $\phi\in[0,2\pi)$, and the polar angle $\theta\in[0,\pi]$. The unit sphere is divided into $N_{\Omega}$ angular bins uniformly in $(\theta,\phi)$.
@@ III.2.1 Identifying Filaments Using the ADPD
Figure 2: Point distribution (upper panel) and heatmap of the ADPD (lower panel) for a system of multiple bars. Each bar is characterized by an orientation angle $\theta_{i}$ relative to the $x$-axis and a finite length $\ell_{i}$.
@@ III.2.2 Analysis lattice’s like structures
Let us move to the case of a regular lattice in 2D. As a reference we consider the case of a random Poisson distribution; in addition we consider shuffled lattices (SL) with different levels of shuffling $\delta$. The SL configurations are generated by randomly displacing each particle from its position on a perfect lattice by a quantity $\delta\leq 1$, expressed in units of the lattice spacing $a$. We refer the interested reader to [12] for a comprehensive analytical and numerical discussion of the two-point correlation function and the power spectrum in such systems.
Figure 3: Angular Distribution of Pairwise Distances (ADPD) for four configurations: a Poisson distribution (top left), a perfect cubic lattice (top right), a shuffled lattice with displacement amplitude $\delta=0.6$ (bottom left), and with $\delta=1.0$ (bottom right).
The ADPD for a random two-dimensional distribution, such as an isotropic Poisson process, is shown in the top-left panel of Figure 3. One observes the that there are no preferred directions, i.e. the ADPD is uniform in angle for any fixed radius $r$. If the ADPD is normalized such that the integral over the interval $\theta\in[0^{\circ},180^{\circ}]$ equals 1, then:
@@
The bottom left panel of Figure 3 shows the ADPD of a shuffled lattice with $\delta=0.6$. Although the coherent lattice structure is partially disrupted, directional peaks remain clearly visible across all pairwise separations. The persistence of narrow features, particularly at $\theta=0^{\circ}$ and $\theta=90^{\circ}$, even with significantly reduced amplitude compared to the perfect lattice, reflects the residual one-dimensional alignment inherited from the original configuration. Relative to the unperturbed case, the peaks are broader and weaker, indicating a partial loss of coherence. Nevertheless, the presence of these anisotropic signatures at both small and large separations suggests that directional ordering persists on large scales despite the imposed perturbations.
A simple $\chi^{2}$ test confirms that the observed peaks both in this case and in the simulations discussed in the following section are statistically significant, demonstrating that the ADPD deviates from a uniform isotropic distribution. As the perturbation amplitude increases ($\delta\gtrsim 0.5$), the peak structure becomes progressively suppressed at all scales. This trend is illustrated in the bottom right panel of Figure 3, where $\delta=1.0$: in this case, the heatmap closely resembles that of a Poisson distribution, with no statistically significant angular features.
@@ III.3 Angular Variance of the ADPD and Comparison with Radial Number Variance
This implies that $\sigma^{2}_{\theta}(r)\sim 1/N^{2}$, and hence becomes negligible in the limit of large $N$. As a result, any excess variance above this Poissonian expectation indicates the presence of anisotropic correlations in the particle distribution. Fig.4 shows the behavior of Eq.12 for some realizations of a Poisson distribution with different number of points.
@@
Figure 4: Angular variance of the Angular Distribution of Pairwise Distances (ADPD) for Poisson distributions with the different number of particles (see labels). A reference line $\propto 1/r$ (see Eq.12) is also shown as a reference.
Figure 5: Angular variance of the ADPD for four different configurations: a single bar, a perfect cubic lattice, a shuffled lattice with displacement amplitude $\delta=0.2$, and with $\delta=0.6$. As reference the angular variance of the ADPD for a Poisson with same number of particles is also reported.
@@ IV Quantifying Local Anisotropy in Cosmological Initial Conditions
The amplitude of these anisotropic modes is directly related to the power spectrum of initial density fluctuations. In linear theory, the variance of the deformation tensor is given by [43]:
@@
where $P(k,z)$ is the power spectrum at redshift $z$. Since the power spectrum evolves as $P(k,z)=D^{2}(z)\,P(k,0)$ in the linear regime, the amplitude of anisotropic deformations grows proportionally to the square of the linear growth factor : $\sigma_{\theta,3D}^{2}(z,r)\propto D^{2}(z).$
@@
where $P(k)$ is the linear matter power spectrum at $z=0$, and $D(z)$ is the linear growth factor. The function $W_{3D}(kr)$ is a smoothing window function that defines the filtering scale: common choices include a Gaussian window function or a Top-hat window function (in real space). Eq.16 expression quantifies how local anisotropies in the deformation tensor depend on the underlying power spectrum and evolve through cosmic time.
Similarly the 2D anisotropy variance can be derived from projecting a 3D power spectrum onto a 2D plane and applying a 2D top-hat window function:
@@
we find
$A(z=0)=3.14\times 10^{5}\;\;(\mathrm{Mpc}/h)^{3}\;.$ (20)
This normalization ensures consistency with the amplitude of density fluctuations inferred from Cosmic Microwave Background observations.
In the range of distance and redshift in which the matter power spectrum linearly evolves with redshift we have
@@
$\Omega_{m}(z)=\frac{\Omega_{m,0}(1+z)^{3}}{E(z)^{2}},\quad\Omega_{\Lambda}(z)=\frac{\Omega_{\Lambda,0}}{E(z)^{2}},$ (25)
@@
$\sigma_{\theta,2D}^{2}(z,r)\sim r^{-4}\;.$ (27)
This is the theoretical behavior to be compared with the measurement of $\sigma^{2}_{\theta}(z,r)$ in the simulations.
@@ V.1.1 Anisotropies in the Initial Conditions: ADPD Analysis
Figure 7: Upper panel: Point distribution of the initial conditions (ICs) for the small-box simulation at redshift $z=9$. Bottom panel: Heatmap of the Angular Distribution of Pairwise Distances (ADPD) corresponding to the same configuration. The heatmap reveals directional anisotropies and preferred orientations in the initial particle distribution.
Figure 7 displays the ADPD heatmap for the initial particle configuration at redshift $z=9$. The distribution is clearly anisotropic, with prominent angular structure observed across a wide range of radial scales. In particular, the range $r\approx 5$$25$ Mpc/h/h exhibits alternating bright and dark bands, corresponding to directional excesses in the pair counts. These indicate the presence of preferred orientations in the underlying point distribution.
Unlike the sharp vertical ridges at $\theta=0^{\circ}$ or $90^{\circ}$ that would be expected from a regular grid or a localized peak for a single bar-like feature, the heatmap reveals curved patterns and broad angular bands. This suggests the presence of multiple filamentary or bar-like structures with different orientations, or possibly a superposition of irregular or diffuse anisotropic components.
Specific angular excesses are observed at distinct angular and radial intervals. For instance, significant enhancements appear in the intervals $\theta\in[10^{\circ},30^{\circ}]$ at $r\sim 20$$30$ Mpc/h/h and $\theta\in[120^{\circ},140^{\circ}]$ at $r\sim 15$ Mpc/h/h. These features are localized in radial extent, implying that the associated structures are finite in size and spatially coherent. Although the amplitude of the anisotropies is modest, their coherence suggests the potential for amplification through gravitational evolution.
The complexity of the observed angular structure supports the interpretation that the initial distribution is inherently anisotropic, comprising multiple substructures with varying orientations and scales. These features do not align with the symmetries of a perfect lattice, but instead reflect a rich and irregular configuration. The heatmap reveals no evidence of global angular symmetry; rather, it appears consistent with a superposition of randomly oriented filaments or bent structures.
@@
Under ideal conditions, CDM initial configurations should satisfy the following: the fluctuation field should be statistically isotropic; the two-point correlation function $\xi(r)$ should decay exponentially, with characteristic scale $r_{c}\sim 15$ Mpc/h/h, and vanish for $r\gg r_{c}$; and the normalized ADPD should be flat, i.e., independent of $\theta$, with constant value $\text{ADPD}(\theta)\sim 1/180^{\circ}$.
In contrast, the ADPD heatmap for this realization shows clear deviations from these expectations. Directional excesses are apparent in $\text{ADPD}(\theta,r)$ across the range $r\lesssim 15$$25$ Mpc/h/h, and the angular distribution contains multiple localized peaks at distinct angles. This indicates a breakdown of isotropy. Furthermore, anisotropic features persist up to $r\sim 30$ Mpc/h/h, well beyond the expected correlation length $r_{c}$, which is inconsistent with the theoretical predictions for a CDM-like initial density field.
@@
we can assess the presence of excess anisotropy. If $\Delta\sigma_{\theta}^{2}(r)>0$, it indicates that the ICs exhibit anisotropies beyond what is expected from Poisson noise alone. In the case of the small box simulation $\sigma^{2}_{\theta,\text{Poisson}}(r)$ is more than 3 orders of magnitude smaller than $\sigma^{2}_{\theta,\text{IC}}(r)$ (see Fig.8). In addition, one can observe that $\sigma^{2}_{\theta,\text{IC}}(r)\approx\text{const.}$, whereas the theoretical expectation scales as $\sim r^{-4}$ (see Eq. 27).
Figure 8: Angular variance of the ADPD computed from the initial conditions (ICs) of the small-box simulation. This diagnostic reveals the presence of directional anisotropies imprinted by the coupling between the isotropic displacement field and the underlying regular lattice.
@@ V.1.2 Time Evolution of Anisotropies in the Small-Box Simulation
Figure 9: Heat-map of the ADPD for or a slice of thickness $\Delta z=5$ Mpc of the small box CDM simulation at different redshifts: first row $z=6.8,5.1,3.7,2.7$, second row $z=1.20.7,0.35,0.05$. The initial conditions are shown in Fig.7.
Figure 10: Point distribution for a the same slices of Fig.9.
Figure 9 displays the evolution of the ADPD heatmap for the small-box simulation at different redshifts. The corresponding projected particle distributions at each redshift are shown in Fig. 10. At the initial redshift $z=9$, the ADPD exhibits sharp peaks at angular positions $\theta=0^{\circ},90^{\circ},180^{\circ},270^{\circ}$ at small separations ($r<2$ Mpc/h/h), inherited from the underlying lattice structure. At larger separations, these peaks broaden and exhibit secondary substructures, revealing more complex angular correlations introduced by the displacement field applied to the lattice.
Figure 11 shows a zoomed view of the 2D point distribution for the initial conditions of the small box at redshifts $z=9$ and $z=6.8$. In the earlier snapshot ($z=9$), the lattice structure is still clearly visible, whereas in the later snapshot ($z=6.8$), it has been almost entirely erased by gravitational evolution. This behavior is consistent with the trends observed in the corresponding ADPD heatmaps.
Figure 11: Left panel: zoom of the point distribution for the IC of the small box at $z=9$. Right panel: zoom of the point distribution at $z=5.1$
As the simulation evolves to lower redshifts ($z<9$), these initial anisotropies persist and gradually amplify. Even after nonlinear clustering becomes dominant at small scales ($r<10$ Mpc/h/h), the ADPD retains the imprint of initial anisotropic features at larger scales. This behavior is consistent with the expectation that gravitational evolution amplifies coherent features present in the initial conditions.
@@
  • At $r\sim 1$ Mpc/h/h, the ADPD evolves rapidly. Peak amplitudes at preferred angles decrease while broadening due to particle motion. The peak structure remains visible until $z\approx 4$, and is effectively frozen by $z\approx 2$, consistent with early nonlinearity at small scales.
@@
Overall, the angular anisotropies present in the initial conditions are amplified by gravitational evolution. This is clearly demonstrated by comparing the ADPD at $z=9$ and $z=0$: although the anisotropic features observed at late times differ in structure, they remain correlated with those imprinted in the initial configuration. This suggests that even small-amplitude directional patterns in the ICs can act as seeds for the filamentary structures that emerge during cosmic evolution, including within the linear and quasi-linear regimes.
@@
where $\sigma^{2}_{\theta}(r,z)$ denotes the angular variance of the ADPD at pair separation $r$ and redshift $z$, and $z_{\mathrm{IC}}$ refers to the redshift at which the initial conditions are set. A value of $G(r,z)>1$ thus reflects the amplification of anisotropic features relative to their initial amplitude. This metric provides a direct way to track the emergence and growth of directional coherence in the matter distribution throughout the simulation. By plotting $G(r,z)$ as a function of $r$, we can interpret the results as follows: If $G(r,z)\gg 1$, this indicates that initial anisotropies have been significantly amplified by gravitational clustering. If $G(r,z)\approx 1$, the angular variance likely originates from Poisson noise and has not been amplified during evolution. As one may note from Fig.12 the variance growth factor is of order $10^{3}$ at $z=0$ showing a very large amplification of initial anisotropies.
Figure 12: Variance growth factor for the small-box simulation, illustrating the amplification of angular anisotropies over time.
We conclude that the anisotropies induced by the coupling between the isotropic displacement field and the underlying regular lattice are not mere numerical noise, but systematic artifacts that persist and grow throughout the simulation. Their presence calls into question the statistical fidelity of lattice-based initial conditions as fair realizations of an isotropic Gaussian field.

### v2

@@ III.1 Correlation function and power spectrum
We will determine the $n$-dimensional correlation function $\xi_{n\text{D}}(r)$ and PS $P_{n\text{D}}(k)$, where $n=2,3$, for the distribution of particles in a simulation box of side length $L$, assuming periodic boundary conditions. For the three-dimensional case ($n=3$), we consider the full volume of the box. We also analyze quasi-two-dimensional slices, i.e., slabs with vertical thickness $\Delta z$ much smaller than the box side length, extending across both the $x$ and $y$ directions. Within each slice, we compute the two-dimensional correlation function $\xi_{2\text{D}}(r)$ and PS $P_{2\text{D}}(k)$. Anisotropies introduced by the lattice used in the pre-ICs are still present even in these two-dimensional slices. Since the numerical analysis in two dimensions is simpler and faster than in three dimensions, we will primarily present results in 2D. We note that, for both simulations, the results presented correspond to a single two-dimensional slice. However, we have repeated the analysis across multiple independent slices, finding similar results in all cases. This confirms that the observed features are robust and not specific to a particular projection. As the statistical properties averaged over spherical (or circular) regions have been extensively studied in the literature, we do not repeat such an analysis in this work.
@@ III.2 Angular Distribution of Pairwise Distances
Estimations of the two-point correlation function or the PS are typically based on the assumption that the underlying distribution is isotropic. However, it is precisely this assumption that we aim to test. To determine whether a point distribution is isotropic or exhibits preferred orientations, one can analyze spatial anisotropies using a range of statistical and geometric methods. We have focused on one particular statistical measure after testing several possibilities. This choice does not imply that it is necessarily the most suitable or optimal diagnostic; rather, it demonstrates that this statistic is effective in identifying anisotropic features in the point distribution.
@@
As mentioned above, given the geometry of the system, we focus on two-dimensional slices with finite thickness. Although structure formation is inherently a fully three-dimensional dynamical process, it is important to consider that particles typically move only a few Mpc over the entire time evolution. Consequently, from a statistical perspective, it is sufficient to isolate a two-dimensional slice of thickness greater than $\sim 5-10$ Mpc in order to track approximately the same subsample of particles throughout the evolution.
Moreover, the symmetry of the lattice used as a pre-ICs justifies reducing the analysis of spatial anisotropies to a two-dimensional slice. This approach is not only numerically more efficient but also retains the essential features of directional correlations that are relevant to the problem.
@@
The generalization of the ADPD to quantify angular anisotropies in a three-dimensional point distribution, is straightforward. The direction of a separation vector is specified by two angles in spherical coordinates: the azimuthal angle $\phi\in[0,2\pi)$, and the polar angle $\theta\in[0,\pi]$. The unit sphere is divided into $N_{\Omega}$ angular bins uniformly in $(\theta,\phi)$.
@@ III.2.1 Identifying Filaments Using the ADPD
Figure 2: Point distribution (upper panel) and heatmap of the Angular Distribution of Pairwise Distances (ADPD, lower panel) for a system composed of multiple bars. Each bar is characterized by an orientation angle $\theta_{i}$ relative to the $x$-axis and a finite length $\ell_{i}$. Length scales are expressed in units of the average distance between nearest neighbors $\Lambda$.
@@ III.2.2 Analysis lattice’s like structures
Let us move to the case of a regular lattice in 2D. As a reference we consider the case of a random Poisson distribution; in addition we consider shuffled lattices (SL) with different levels of shuffling $\delta$. The SL configurations are generated by randomly displacing each particle from its position on a perfect lattice by a quantity $\delta\leq 1$, expressed in units of the lattice spacing $a$. We refer the interested reader to [12] for a comprehensive analytical and numerical discussion of the two-point correlation function and the PS in such systems.
Figure 3: Angular Distribution of Pairwise Distances (ADPD) for four configurations: a Poisson distribution (top left), a perfect cubic lattice (top right), a shuffled lattice with displacement amplitude $\delta=0.6$ (bottom left), and with $\delta=1.0$ (bottom right). For improved visual interpretation, the color scale in each panel has been individually adjusted to enhance contrast and highlight relevant structures in the ADPD. Length scales are expressed in units of the average distance between nearest neighbors $\Lambda$.
The ADPD for a random two-dimensional distribution, such as an isotropic Poisson process, is shown in the top-left panel of Fig. 3. One observes the that there are no preferred directions, i.e. the ADPD is uniform in angle for any fixed radius $r$. If the ADPD is normalized such that the integral over the interval $\theta\in[0^{\circ},180^{\circ}]$ equals 1, then:
@@
The bottom left panel of Fig. 3 shows the ADPD of a shuffled lattice with $\delta=0.6$. Although the coherent lattice structure is partially disrupted, directional peaks remain clearly visible across all pairwise separations. The persistence of narrow features, particularly at $\theta=0^{\circ}$ and $\theta=90^{\circ}$, even with significantly reduced amplitude compared to the perfect lattice, reflects the residual one-dimensional alignment inherited from the original configuration. Relative to the unperturbed case, the peaks are broader and weaker, indicating a partial loss of coherence. Nevertheless, the presence of these anisotropic signatures at both small and large separations suggests that directional ordering persists on large scales despite the imposed perturbations.
A simple $\chi^{2}$ test confirms that the observed peaks both in this case and in the simulations discussed in the following section are statistically significant, demonstrating that the ADPD deviates from a uniform isotropic distribution. As the perturbation amplitude increases ($\delta\gtrsim 0.5$), the peak structure becomes progressively suppressed at all scales. This trend is illustrated in the bottom right panel of Fig. 3, where $\delta=1.0$: in this case, the heatmap closely resembles that of a Poisson distribution, with no statistically significant angular features.
The key message of Fig. 3 can be summarized as follows: a lattice or a distribution close to a lattice produces regular, anisotropic patterns in the ADPD heatmap, whereas a random (Poisson) distribution does not. In particular, the top-left panel of Fig. 3 is clearly one of the noisiest, as it corresponds to a Poisson distribution and is thus dominated by random fluctuations. The bottom-right panel, showing a shuffled lattice with $\delta=1.0$, also closely resembles a Poisson distribution, as expected.
@@ III.3 Angular Variance of the ADPD and Comparison with Radial Number Variance
This implies that $\sigma^{2}_{\theta}(r)\propto 1/N^{2}$, and hence becomes negligible in the limit of large $N$. As a result, any excess variance above this Poissonian expectation indicates the presence of anisotropic correlations in the particle distribution. Fig.4 shows the behavior of Eq.12 for some realizations of a Poisson distribution with different number of points.
@@
Figure 4: Angular variance of the Angular Distribution of Pairwise Distances for Poisson distributions in a box of side $L=400$ Mpc/$h$ with the different number of particles (see labels). A line $\propto 1/r$ (see Eq.12) is also shown as a reference.
Figure 5: Angular variance of the Angular Distribution of Pairwise Distances for four different configurations: a single bar, a perfect cubic lattice, a shuffled lattice with displacement amplitude $\delta=0.2$, and with $\delta=0.6$. All distributions are generated in a box of side $L=400$ Mpc/$h$ and have the same number of points $N=10^{5}$. As reference the angular variance of the ADPD for a Poisson with same number of particles is also reported.
@@ IV Quantifying Local Anisotropy in Cosmological Initial Conditions
The amplitude of these anisotropic modes is directly related to the PS of initial density fluctuations. In linear theory, the variance of the deformation tensor is given by [47]:
@@
where $P(k,z)$ is the PS at redshift $z$. Since the PS evolves as $P(k,z)=D^{2}(z)\,P(k,0)$ in the linear regime, the amplitude of anisotropic deformations grows proportionally to the square of the linear growth factor : $\sigma_{\theta,3D}^{2}(z,r)\propto D^{2}(z).$
@@
where $P(k)$ is the linear matter PS at $z=0$, and $D(z)$ is the linear growth factor. The function $W_{3D}(kr)$ is a smoothing window function that defines the filtering scale: common choices include a Gaussian window function or a Top-hat window function (in real space). Eq.16 expression quantifies how local anisotropies in the deformation tensor depend on the underlying PS and evolve through cosmic time.
Similarly the 2D anisotropy variance can be derived from projecting a 3D PS onto a 2D plane and applying a 2D top-hat window function:
@@
where $R_{8}=R=8\,h^{-1}\mathrm{Mpc}$, we find
$A(z=0)=\frac{\sigma_{8}^{2}}{\frac{1}{2\pi^{2}}\int_{0}^{\infty}k^{2}\hat{P}(k)\,W_{\rm TH}^{2}(kR_{8})\,dk}$  (20)
$=3.14\times 10^{5}\;\;(\mathrm{Mpc}/h)^{3}\;,$
where $\hat{P}(k)$ is the shape function. This normalization ensures consistency with the amplitude of density fluctuations inferred from Cosmic Microwave Background observations.
In the range of distance and redshift in which the matter PS linearly evolves with redshift we have
@@
$\Omega_{m}(z)=\frac{\Omega_{m,0}(1+z)^{3}}{E(z)^{2}}\;,\;\;\Omega_{\Lambda}(z)=\frac{\Omega_{\Lambda,0}}{E(z)^{2}},$ (25)
@@
$\sigma_{\theta,2D}^{2}(z,r)\propto r^{-4}\;.$ (27)
This behaviour provides, in principle, a natural reference for comparison with measurements of $\sigma^{2}_{\theta}(z,r)$ in simulations. However, Eq. 27 applies to a continuous field, whereas the angular variance measured in a discrete point distribution depends on the specific discretisation adopted, and its amplitude cannot be predicted a priori. By contrast, as long as the point distribution remains uniform and statistically isotropic, the asymptotic scaling of $\sigma^{2}_{\theta}(r)$ — analogous to that of a Poisson distribution — and the corresponding featureless appearance of the ADPD heatmap constitute robust, model-independent expectations.
@@ V Results in simulations
The substantial differences in particle number (i.e., resolution) and box size between the two simulations allow us to investigate the impact of finite-size effects on statistical measures such as the two-point correlation function and the ADPD. In both box cases, we select a two-dimensional thin slice of side length $L$ and thickness $\Delta z=5$ Mpc/h/h for the small box and $\Delta z=10$ Mpc/h/h for the large box.
@@ V.1.1 Anisotropies in the Initial Conditions: ADPD Analysis
[…]
@@
Figure 7: Upper panel: Point distribution in a zoomed region of the initial conditions for the small-box simulation at redshift $z_{\mathrm{IC}}=49$. (units are in Mpc/$h$). Bottom panel: Heatmap of the Angular Distribution of Pairwise Distances (ADPD) corresponding to the same configuration. Length scales are expressed in units of the average distance between nearest neighbors $\Lambda$.
Under ideal conditions, CDM initial configurations should satisfy the following: the fluctuation field should be statistically isotropic; the two-point correlation function $\xi(r)$ should decay exponentially, with characteristic scale $r_{c}\sim 15$ Mpc/h/h, and vanish for $r\gg r_{c}$; and the normalized ADPD should be flat, i.e., independent of $\theta$, with constant value $\text{ADPD}(\theta)\sim 1/180^{\circ}$. In contrast, the ADPD heatmap for this realization shows clear deviations from these expectations.
@@
we can assess the presence of excess anisotropy. If $\Delta\sigma_{\theta}^{2}(r)>0$, it indicates that the ICs exhibit anisotropies beyond what is expected from Poisson noise alone. In the case of the small box simulation $\sigma^{2}_{\theta,\text{Poisson}}(r)$ is more than 5 orders of magnitude smaller than $\sigma^{2}_{\theta,\text{IC}}(r)$ (see Fig.8). In addition, one can observe that, at large enough $r$, $\sigma^{2}_{\theta,\text{IC}}(r)\approx\text{const.}$, whereas the theoretical expectation scales as $\propto r^{-4}$ (see Eq. 27).
Figure 8: Angular variance of the ADPD computed from the initial conditions of the small-box simulation. This diagnostic reveals the presence of directional anisotropies imprinted by the coupling between the isotropic displacement field and the underlying regular lattice. Length scales are expressed in units of the average distance between nearest neighbors $\Lambda$.
@@ V.1.2 Time Evolution of Anisotropies in the Small-Box Simulation
Figure 9: Heat-map of the ADPD for or a slice of thickness $\Delta z=5$ Mpc of the small box CDM simulation at different redshifts: first row $z=9.0,5.1,3.7,2.7$, second row $z=1.2,0.7,0.35,0.05$. The ICs are shown in Fig.7. Length scales are expressed in units of the average distance between nearest neighbors $\Lambda$.
Figure 10: Point distribution for a the same slices of Fig.9. Length scales are expressed in units of the average distance between nearest neighbors $\Lambda$.
Fig. 9 displays the evolution of the ADPD heatmap for the small-box simulation at different redshifts. The corresponding projected particle distributions at each redshift are shown in Fig. 10. At redshift $z=9$, the heatmap of the ADPD still exhibits distinct patterns, notably sharp peaks at angular positions $\theta=0^{\circ},90^{\circ},180^{\circ},270^{\circ}$ for small separations ($r<2$ Mpc/h/h), reflecting the underlying lattice structure (see Fig.7). At larger separations, these peaks become broader and display secondary substructures, indicating the emergence of more complex angular correlations induced by the displacement field applied to the initial lattice. These features remain visible in the ADPD heatmap at all subsequent redshifts, with their overall amplitude progressively increasing.
Fig. 11 shows a zoomed view of the two-dimensional point distribution for the small-box simulation at redshifts $z=9$ and $z=5.1$. In the earlier snapshot ($z=9$), the lattice structure—clearly visible in the ICs (see Fig. 7)—is still apparent, whereas in the later snapshot ($z=5.1$) it has been partially erased by gravitational evolution. This evolution is consistent with the trends observed in the corresponding ADPD heatmaps.
Figure 11: Left panel: zoom of the point distribution for the IC of the small box at $z=9$. Right panel: zoom of the point distribution at $z=5.1$. Length scales are expressed in units of the average distance between nearest neighbors $\Lambda$.
As the simulation evolves to lower redshifts ($z<9$), these initial anisotropies persist and gradually amplify. Even after nonlinear clustering becomes dominant at small scales ($r<10$ Mpc/h/h), the ADPD retains the imprint of initial anisotropic features at larger scales. This behavior is consistent with the expectation that gravitational evolution amplifies coherent features present in the ICs.
@@
  • At $r\sim 10$ Mpc/h/h, the ADPD evolves rapidly. Peak amplitudes at preferred angles decrease while broadening due to particle motion. The peak structure remains visible until $z\approx 4$, and is effectively frozen by $z\approx 2$, consistent with early nonlinearity at small scales.
@@
Overall, the angular anisotropies present in the ICs are amplified by gravitational evolution. This is clearly demonstrated by comparing the ADPD at $z=9$ and $z=0$: although the anisotropic features observed at late times differ in structure, they remain correlated with those imprinted in the initial configuration. This suggests that even small-amplitude directional patterns in the ICs can act as seeds for the filamentary structures that emerge during cosmic evolution, including within the linear and quasi-linear regimes.
@@
where $\sigma^{2}_{\theta}(r,z)$ denotes the angular variance of the ADPD at pair separation $r$ and redshift $z$, and $z_{\mathrm{IC}}$ refers to the redshift at which the ICs are set. A value of $G(r,z)>1$ thus reflects the amplification of anisotropic features relative to their initial amplitude. This metric provides a direct way to track the emergence and growth of directional coherence in the matter distribution throughout the simulation. By plotting $G(r,z)$ as a function of $r$, we can interpret the results as follows: If $G(r,z)\gg 1$, this indicates that initial anisotropies have been significantly amplified by gravitational clustering. If $G(r,z)\approx 1$, the angular variance likely originates from Poisson noise and has not been amplified during evolution. As one may note from Fig.12 the variance growth factor is of order $10^{2}$ at $z=0$ showing a very large amplification of initial anisotropies.
Figure 12: Variance growth factor for the small-box simulation, illustrating the amplification of angular anisotropies over time. Length scales are expressed in units of the average distance between nearest neighbors $\Lambda$.
We conclude that the anisotropies induced by the coupling between the isotropic displacement field and the underlying regular lattice are not mere numerical noise, but systematic artifacts that persist and grow throughout the simulation. Their presence calls into question the statistical fidelity of lattice-based ICs as fair realizations of an isotropic Gaussian field.